# 3D Diffuser Actor: Multi-task 3D Robot Manipulation with Iterative Error Feedback

### v2

@@ Abstract
We marry diffusion policies and 3D scene representations for robot manipulation. Diffusion policies learn the action distribution conditioned on the robot and environment state using conditional diffusion models. They have recently shown to outperform both deterministic and alternative state-conditioned action distribution learning methods. 3D robot policies use 3D scene feature representations aggregated from a single or multiple camera views using sensed depth. They have shown to generalize better than their 2D counterparts across camera viewpoints. We unify these two lines of work and present 3D Diffuser Actor, a neural policy architecture that, given a language instruction, builds a 3D representation of the visual scene and conditions on it to iteratively denoise 3D rotations and translations for the robot’s end-effector. At each denoising iteration, our model represents end-effector pose estimates as 3D scene tokens and predicts the 3D translation and rotation error for each of them, by featurizing them using 3D relative attention to other 3D visual and language tokens. 3D Diffuser Actor sets a new state-of-the-art on RLBench with an absolute performance gain of 18.1% over the current SOTA on a multi-view setup and an absolute gain of 13.1% on a single-view setup. On the CALVIN benchmark, it outperforms the current SOTA in the setting of zero-shot unseen scene generalization by being able to successfully run 0.2 more tasks, a 7% relative increase. It also works in the real world from a handful of demonstrations. We ablate our model’s architectural design choices, such as 3D scene featurization and 3D relative attentions, and show they all help generalization. Our results suggest that 3D scene representations and powerful generative modeling are keys to efficient robot learning from demonstrations.
footnotetext: ${}^{\dagger}$qual contribution
I Introduction
Figure 1: 3D Diffuser Actor sets a new state-of-the-art on RLBench [32] on a multi-view setup and on CALVIN [49] benchmarks on a zero-shot long-horizon setup. On RLBench, we train and evaluate on the PerAct setup. On CALVIN, we train on environments A, B, C and test on environment D. 3D Diffuser Actor achieves state-of-the-art performance on both benchmarks.
Figure 2: 3D Diffuser Actor is a language and vision conditioned action diffusion policy that learns to predict end-effector keyposes and trajectories from demonstrations. (a) An image encoder extracts a 2D image feature map from each camera view; each feature token corresponds to a fixed-size patch in the input image. (b) We lift 2D feature tokens to 3D using sensed depth. (c) A 3D relative position attention Transformer takes as input the 3D scene tokens, an instruction text encoding, a short history of end-effector poses and the current noisy estimate of the end-effector future trajectory (or keypose) and outputs the noise in the form of 3D translations and rotations.
@@ I Introduction
Many robot manipulation tasks are inherently multimodal: at any point during task execution, there may be multiple actions which yield task-optimal behavior. Indeed, human demonstrations often contain diverse ways that a task can be accomplished. A natural choice is then to treat policy learning as a distribution learning problem: instead of representing a policy as a deterministic map $\pi_{\theta}(x)$, learn the entire distribution of actions conditioned on the current robot state $p(y|x)$ [27, 76, 25, 66]. Recent works use diffusion models for learning such state-conditioned action distributions for robot manipulation policies from demonstrations [54, 11, 60] and show they outperform deterministic or other alternatives, such as variational autoencoders [47], mixture of Gaussians [10], combination of classification and regression objectives [66], or energy-based objectives [18]. Specifically, they exhibit better action distribution coverage and higher fidelity (less mode hallucination) than alternative formulations [54]. They have so far been used with low-dimensional engineered state representations [54] or 2D image encodings [11].
Lifting features from perspective views to a bird’s eye view (BEV) or 3D robot workspace map has shown strong results in robot learning from demonstrations [69, 33, 22, 19]. Policies that use such 2D-to-BEV or 2D-to-3D scene encodings generalize better than their 2D counterparts across camera viewpoints and can handle novel camera viewpoints at test time [19, 22]. Our conjecture is that this improved performance comes from the fact that the 3D visual scene content and the robot’s end-effector poses live in a common 3D space. BEV or 3D policy formulations e.g., Transporter Networks [84, 29], C2F-ARM [33], PerAct [69], Act3D [19] and Robot View Transformer [22], discretize the robot’s workspace for localizing the robot’s end-effector, such as the 2D BEV map for pick and place actions [84], the 3D workspace [19], or multiview re-projected images [22]. Such 3D policies have not been combined yet with diffusion objectives.
@@
In this paper, we propose 3D Diffuser Actor, a model that marries diffusion policies for handling action multimodality and 3D scene encodings for effective spatial reasoning. 3D Diffuser Actor trains a denoising neural network that takes as input a 3D scene visual encoding, the current estimate of the end-effector’s future trajectory (3D location and orientation), as well as the diffusion iteration index, and predicts the error in 3D translation and rotation. Our model achieves translation equivariance in prediction by representing the current estimate of the robot’s end-effector trajectory as 3D scene tokens and featurizing them jointly with the visual tokens using relative-position 3D attentions [67, 74], as shown in Figure 2.
@@
We test 3D Diffuser Actor in learning robot manipulation policies from demonstrations on the simulation benchmarks of RLBench [32] and CALVIN [49], as well as in the real world. Our model sets a new state-of-the-art on RLBench, outperforming existing 3D policies and 2D diffusion policies with a 18.1% absolute gain on multi-view setups and 13.1% on single-view setups (Figure 1). On CALVIN, our model outperforms the current SOTA in the setting of zero-shot unseen scene generalization by a 7% relative gain (Figure 1). We additionally show that 3D Diffuser Actor outperforms all existing policy formulations that either do not use 3D scene representations or do not use action diffusion, as well as ablative versions of our model that do not use 3D relative attentions. We further show 3D Diffuser Actor can learn multi-task manipulation in the real world across 12 tasks from a handful of real world demonstrations. Simulation and real-world robot execution videos, as well as our code and trained checkpoints are publicly available on our website: 3d-diffuser-actor.github.io.
II Related Work
@@
Learning robot manipulation policies from demonstrations Though state-to-action mappings are typically multimodal, earlier works on learning from demonstrations train deterministic policies with behavior cloning [56, 3]. To better handle action multimodality, other approaches discretize action dimensions and use cross entropy losses [45, 69, 84]. However, the number of bins needed to approximate a continuous action space grows exponentially with increasing dimensionality. Generative adversarial networks [27, 76, 16], variational autoencoders [47] and combined Categorical and Gaussian distributions [66, 23, 14] have been used to learn from multimodal demonstrations. Nevertheless, these models tend to be sensitive to hyperparameters, such as the number of clusters used [66]. Implicit behaviour cloning represents distributions over actions by using Energy-Based Models (EBMs) [18, 75]. Optimization in EBMs amounts to searching the energy landscape for minimal-energy action, given a state. EBMs are inherently multimodal, since the learned landscape can have more than one minima. EBMs are also composable, which makes them suitable for combining action distributions with additional constraints during inference [20]. Diffusion models [72, 28] are a powerful class of generative models related to EBMs in that they model the score of the distribution, else, the gradient of the energy, as opposed to the energy itself [71, 63]. The key idea behind diffusion models is to iteratively transform a simple prior distribution into a target distribution by applying a sequential denoising process. They have been used for modeling state-conditioned action distributions in imitation learning [62, 77, 54, 78, 60, 51] from low-dimensional input, as well as from visual sensory input, and show both better mode coverage and higher fidelity in action prediction than alternatives. They have not been yet combined with 3D scene representations.
@@
Diffusion models in robotics Beyond policy representations in imitation learning, diffusion models have been used to model cross-object and object-part arrangements [43, 70, 51, 17, 20] and visual image subgoals [36, 15, 1, 2]. They have also been used successfully in offline reinforcement learning [7, 82, 24], where they model the state-conditioned action trajectory distribution [24, 7] or state-action trajectory distribution [35, 26, 79]. ChainedDiffuser [81] proposes to replace motion planners, commonly used for keypose to keypose linking, with a trajectory diffusion model that conditions on the 3D scene feature cloud and the predicted target 3D keypose to denoise a trajectory from the current to the target keypose. 3D Diffuser Actor instead predicts the next 3D keypose for the robot’s end-effector alongside the linking trajectory, which is a much harder task than linking two given keyposes. Lastly, image diffusion models have been used for augmenting the conditioning images input to robot policies to help the latter generalize better [40, 9, 46].
@@
2D and 3D scene representations for robot manipulation End-to-end image-to-action policy models, such as RT-1 [4], RT-2[5], GATO [59], BC-Z [34], RT-X [53], Octo [52] and InstructRL [42] leverage transformer architectures for the direct prediction of 6-DoF end-effector poses from 2D video input. However, this approach comes at the cost of requiring thousands of demonstrations to implicitly model 3D geometry and adapt to variations in the training domains. Another line of research is centered around Transporter networks [84, 65, 68, 20], demonstrating remarkable few-shot generalization by framing end-effector pose prediction as pixel classification. Nevertheless, these models are usually confined to top-down 2D planar environments with simple pick-and-place primitives. Direct extensions to 3D, exemplified by C2F-ARM [33] and PerAct [69], involve voxelizing the robot’s workspace and learning to identify the 3D voxel containing the next end-effector keypose. However, this becomes computationally expensive as resolution requirements increase. Consequently, related approaches resort to either coarse-to-fine voxelization, equivariant networks [29] or efficient attention operations [30] to mitigate computational costs. Act3D [19] foregoes 3D scene voxelization altogether; it instead computes a 3D action map of variable spatial resolution by sampling 3D points in the empty workspace and featurizing them using cross-attentions to the 3D physical points. Robotic View Transformer (RVT) [22] re-projects the input RGB-D image to alternative image views, featurizes those and lifts the predictions to 3D to infer 3D locations for the robot’s end-effector. Both Act3D and RVT show currently the highest performance on RLBench [32]. Our model outperforms them by a big margin, as we show in the experimental section.
Instruction-conditioned long-horizon policies To generate actions based on language instructions, early works [44, 48, 4] employ a text encoder to map language instructions to latent features, then they deploy 2D policies that condition on these language latents to predict actions. HULC++ [50] alternates between a free-space policy that reaches subgoals (next object to interact with) and a local policy [48] for low-level interaction. However, these models struggle to infer 3D actions precisely from 2D observations and language instructions and do not generalize to new environments with different texture. Recent works [2, 38, 80] explore the use of large-scale pre-training to boost their performance. SuSIE [2] proposes to synthesize visual subgoals using InstructPix2Pix [6], an instruction-conditioned image generative model pre-trained on a subset of 5 billion images [64]. Actions are then predicted by a low-level diffusion policy conditioned on both current observations and synthesized visual goals. RoboFlamingo [38] fine-tunes existing vision-language models, pre-trained for solving vision and language tasks, to predict the robot’s actions. GR-1 [80] pre-trains a GPT-style transformer for video generation on a massive video corpus, then jointly trains the model for predicting robot actions and future observations. We show that 3D Diffuser Actor can exceed the performance of these policies thanks to effective use of 3D representations.
Figure 3: 3D Diffuser Actor architecture. Top: 3D Diffuser Actor is a denoising diffusion probabilistic model of the robot 3D trajectory conditioned on sensory input, language goals and proprioceptive information (action history). The model is a 3D relative position denoising transformer that featurizes jointly the scene and the current noisy estimate for robot’s future action trajectory through 3D relative-position attentions. 3D Diffuser Actor outputs position and rotation residuals for denoising, as well as the end-effector’s state (open/close). Bottom: During inference, 3D Diffuser Actor iteratively denoises the current estimate for the robot’s future trajectory, where the initial estimate is initialized from pure noise.
III Method
The architecture of 3D Diffuser Actor is shown in Figure 3. It is a conditional diffusion model that takes as input visual observations, a language instruction, a short history of the robot’s end-effector’s pose and the current estimate for the robot’s future action trajectory, and predicts the error in the end-effector’s 3D translations and 3D orientations for each predicted timestep. We review Denoising Diffusion Probabilistic models in Section III-A and describe the architecture and training details of our model in Section III-B.
III-A Denoising Diffusion Probabilistic Models
A diffusion model learns to model a probability distribution $p(x)$ by inverting a process that gradually adds noise to a sample $x$. For us, $x$ represents a sequence of 3D translations and 3D rotations for the robot’s end-effector. The diffusion process is associated with a variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$, which defines how much noise is added at each time step. The noisy version of sample $x$ at time $t$ can then be written as $x_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1}),$ is a sample from a Gaussian distribution (with the same dimensionality as $x$), $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. The denoising process is modeled by a neural network $\hat{\epsilon}=\epsilon_{\theta}(x_{t};t)$ that takes as input the noisy sample $x_{t}$ and the noise level $t$ and tries to predict the noise component $\epsilon$.
Diffusion models can be easily extended to draw samples from a distribution $p(x|\textbf{c})$ conditioned on input c, which is added as input to the network $\epsilon_{\theta}$. For us c is the visual scene captured by one or more calibrated RGB-D images, a language instruction, as well as a short history of the robot’s end-effector’s poses. Given a collection of $\mathcal{D}=\{(x^{i},\textbf{c}^{i})\}_{i=1}^{N}$ of end-effector trajectories $x^{i}$ paired with observation and robot history context $\textbf{c}^{i}$, the denoising objective becomes:
$\mathcal{L}_{\text{diff}}(\theta;\mathcal{D})=\tfrac{1}{|\mathcal{D}|}\sum_{x^{i},\textbf{c}^{i}\in\mathcal{D}}||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x^{i}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\textbf{c}^{i},t)-\epsilon||.$ (1)
This loss corresponds to a reweighted form of the variational lower bound for $\log p(x|\textbf{c})$ [28].
In order to draw a sample from the learned distribution $p_{\theta}(x|\textbf{c})$, we start by drawing a sample $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$. Then, we progressively denoise the sample by iterated application of $\epsilon_{\theta}$ $T$ times according to a specified sampling schedule [28, 73], which terminates with $x_{0}$ sampled from $p_{\theta}(x)$:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t,\textbf{c})\right)+\frac{1-\bar{\alpha}_{t+1}}{1-\bar{\alpha}_{t}}\beta_{t}\mathbf{z},$ (2)
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
III-B 3D Diffuser Actor
We next describe the architecture of our model that predicts the end-effector’s trajectory error given the current noisy estimate, a visual scene encoding, a language instruction, the denoising iteration index and the end-effector’s pose history.

### v3

# Tsung-Wei Ke   Nikolaos Gkanatsios    Katerina Fragkiadaki

@@ Abstract
Diffusion policies are conditional diffusion models that learn robot action distributions conditioned on the robot and environment state. They have recently shown to outperform both deterministic and alternative action distribution learning formulations. 3D robot policies use 3D scene feature representations aggregated from a single or multiple camera views using sensed depth. They have shown to generalize better than their 2D counterparts across camera viewpoints. We unify these two lines of work and present 3D Diffuser Actor, a neural policy equipped with a novel 3D denoising transformer that fuses information from the 3D visual scene, a language instruction and proprioception to predict the noise in noised 3D robot pose trajectories. 3D Diffuser Actor sets a new state-of-the-art on RLBench with an absolute performance gain of 18.1% over the current SOTA on a multi-view setup and an absolute gain of 13.1% on a single-view setup. On the CALVIN benchmark, it improves over the current SOTA by a 9% relative increase. It also learns to control a robot manipulator in the real world from a handful of demonstrations. Through thorough comparisons with the current SOTA policies and ablations of our model, we show 3D Diffuser Actor’s design choices dramatically outperform 2D representations, regression and classification objectives, absolute attentions, and holistic non-tokenized 3D scene embeddings.
[…]
1 Introduction
@@
Many robot manipulation tasks are inherently multimodal: at any point during task execution, there may be multiple actions which yield task-optimal behavior. Indeed, human demonstrations often contain diverse ways that a task can be accomplished. A natural choice is then to treat policy learning as a distribution learning problem: instead of representing a policy as a deterministic map $\pi_{\theta}(x)$, learn the entire distribution of actions conditioned on the current robot state $p(y|x)$ [DBLP:journals/corr/HoE16, tsurumine2022goalaware, DBLP:journals/corr/HausmanCSSL17, shafiullah2022behavior].
Recent works use diffusion objectives for learning such state-conditioned action distributions for robot manipulation policies from demonstrations [pearce2023imitating, chi2023diffusion, reuss2023goal]. They outperform deterministic or other alternatives, such as variational autoencoders [IRIS], mixture of Gaussians [10.1145/1329125.1329407], combination of classification and regression objectives [shafiullah2022behavior], or energy-based objectives [DBLP:journals/corr/abs-2109-00137]. They typically use either low-dimensional (oracle) states [pearce2023imitating] or 2D images [chi2023diffusion] as their scene representation.
At the same time, 3D robot policies build scene representations by “lifting" features from perspective views to a 3D robot workspace based on sensed depth and camera extrinsics [zeng2021transporter, huang2024fourier, james2022coarse, shridhar2023perceiver, gervet2023act3d, goyal2023rvt]. They have shown to generalize better than 2D robot policies across camera viewpoints and to handle novel camera viewpoints at test time [gervet2023act3d, goyal2023rvt]. We conjecture this improved performance comes from the fact that the visual scene tokens and the robot’s actions interact in a common 3D space, that is robust to camera viewpoints, while in 2D policies the neural network need to learn the 2D-to-3D mapping implicitly.
@@
In this work, we marry diffusion for handling multimodality in action prediction with 3D scene representations for effective spatial reasoning. We propose 3D Diffuser Actor, a novel 3D denoising policy transformer that takes as input a tokenized 3D scene representation, a language instruction and a noised end-effector’s future translation and rotation trajectory, and predicts the error in translations and rotations for the robot’s end-effector. The model represents both the scene tokens and the end-effector locations in the same 3D space and fuses them with relative-position 3D attentions [shaw2018selfattention, su2021roformer], which achieves translation equivariance and helps generalization.
@@
We test 3D Diffuser Actor in learning robot manipulation policies from demonstrations on the simulation benchmarks of RLBench [james2020rlbench] and CALVIN [mees2022calvin], as well as in the real world. 3D Diffuser Actor sets a new state-of-the-art on RLBench with a 18.1% absolute gain on multi-view setups and 13.1% on single-view setups, outperforming existing 3D policies and 2D diffusion policies. On CALVIN, it outperforms the current SOTA in the setting of zero-shot unseen scene generalization by a 9% relative gain. We further show 3D Diffuser Actor can learn multi-task manipulation in the real world across 12 tasks from a handful of real-world demonstrations. We empirically show that 3D Diffuser Actor outperforms all existing policy formulations, which either do not use 3D scene representations, or do not use action diffusion. We further compare against ablative versions of our model and show the importance of the 3D relative attentions.
Our contributions: The main contribution of this work is to combine 3D scene representations and diffusion objectives for learning robot policies from demonstrations. 3D robot policies have not yet been combined with diffusion objectives. An exception is ChainedDiffuser [xian2023chaineddiffuser], that uses diffusion models as a drop-in replacement for motion planners, rather than manipulation policies, since it relies on other learning-based policies (Act3D [gervet2023act3d]) to supply the target 3D keypose to reach. We compare against ChainedDiffuser in our experiments and show we greatly outperform it.
Concurrent work: Concurrent to our effort, 3D diffusion policy [Ze2024DP3] shares a similar goal of combining 3D representations with diffusion objectives for learning manipulation from demonstrations. Though the two works share the same goal, they have very different architectures. Unlike 3D Diffuser Actor, the model of [Ze2024DP3] does not condition on a tokenized 3D scene representation but rather on a holistic 1D embedding pooled from the 3D scene point cloud. We compare 3D Diffuser Actor against [Ze2024DP3] in our experiments and show it greatly outperforms it. We believe this is because tokenized scene representations, used in 3D Diffuser Actor, are robust to scene changes: if a part of the scene changes, only the corresponding subset of 3D scene tokens is affected. In contrast, holistic scene embeddings pooled across the scene are always affected by any scene change. Thanks to this spatial disentanglement of the 3D scene tokenization, 3D Diffuser Actor generalizes better.
Our models and code will be publicly available upon publication. Videos of our manipulation results are available at https://sites.google.com/view/3d-diffuser-actor.
2 Related Work
@@
Learning manipulation policies from demonstrations Earlier works on learning from demonstrations train deterministic policies with behavior cloning [Pomerleau-1989-15721, DBLP:journals/corr/BojarskiTDFFGJM16]. To better handle action multimodality, approaches discretize action dimensions and use cross entropy losses [semanetlang, shridhar2023perceiver, zeng2021transporter]. Generative adversarial networks [DBLP:journals/corr/HoE16, tsurumine2022goalaware, NEURIPS2019_c8d3a760], variational autoencoders [IRIS], combined Categorical and Gaussian distributions [shafiullah2022behavior, guhur2023instruction, Cui2022FromPT] and Energy-Based Models (EBMs) [DBLP:journals/corr/abs-2109-00137, ta2022conditional, gkanatsios2023energy] have been used to learn from multimodal demonstrations. Diffusion models [sohldickstein2015deep, DDPM] are a powerful class of generative models related to EBMs in that they model the score of the distribution, else, the gradient of the energy, as opposed to the energy itself [singh2023revisiting, salimans2021should]. The key idea behind diffusion models is to iteratively transform a simple prior distribution into a target distribution by applying a sequential denoising process. They have been used for modeling state-conditioned action distributions in imitation learning [ryu2023diffusion, urain2023se, pearce2023imitating, wang2022diffusion, reuss2023goal, mishra2023reorientdiff] from low-dimensional input, as well as from visual sensory input, and show both better mode coverage and higher fidelity in action prediction than alternatives.
@@
Diffusion models in robotics Beyond policy representations in imitation learning, diffusion models have been used to model cross-object and object-part arrangements [liu2022structdiffusion, simeonov2023shelving, mishra2023reorientdiff, fang2023dimsam, gkanatsios2023energy], visual image subgoals [kapelyukh2023dall, dai2023learning, ajay2023hip, black2023zero], and offline reinforcement learning [chen2023offline, yang2024diffusiones, hansen2023idql]. ChainedDiffuser [xian2023chaineddiffuser] proposes to replace motion planners, commonly used for keypose to keypose linking, with a trajectory diffusion model that conditions on the 3D scene feature cloud and the predicted target 3D keypose to denoise a trajectory from the current to the target keypose. It uses a diffusion model that takes as input 3D end-effector keyposes predicted by Act3D [gervet2023act3d] and a 3D representation of the scene to infer robot end-effector trajectories that link the current end-effector pose to the predicted one. 3D Diffuser Actor instead predicts the next 3D keypose for the robot’s end-effector alongside the linking trajectory, which is a much harder task than linking two given keyposes. 3D Diffusion Policy [Ze2024DP3] also combines 3D scene representations with diffusion objectives but uses 1D point cloud embeddings. We compare against both ChainedDiffuser and 3D Diffusion Policy in our experimental section and show we greatly outperform them.
@@
2D and 3D scene representations for robot manipulation End-to-end image-to-action policy models, such as RT-1 [brohan2022rt], RT-2 [brohan2023rt2], GATO [reed2022generalist], BC-Z [jang2022bc], RT-X [Padalkar2023OpenXR], Octo [octo_2023] and InstructRL [liu2022instruction] leverage transformer architectures for the direct prediction of 6-DoF end-effector poses from 2D video input. However, this approach comes at the cost of requiring thousands of demonstrations to implicitly model 3D geometry and adapt to variations in the training domains. 3D scene-to-action policies, exemplified by C2F-ARM [james2022coarse] and PerAct [shridhar2023perceiver], involve voxelizing the robot’s workspace and learning to identify the 3D voxel containing the next end-effector keypose. However, this becomes computationally expensive as resolution requirements increase. Consequently, related approaches resort to either coarse-to-fine voxelization, equivariant networks [huang2024fourier] or efficient attention operations [jaegle2021perceiver] to mitigate computational costs. Act3D [gervet2023act3d] foregoes 3D scene voxelization altogether; it instead computes a 3D action map of variable spatial resolution by sampling 3D points in the empty workspace and featurizing them using cross-attentions to the 3D physical scene points. Robotic View Transformer (RVT) [goyal2023rvt] re-projects the input RGB-D image to alternative image views, featurizes those and lifts the predictions to 3D to infer 3D locations for the robot’s end-effector.
3D Diffuser Actor builds upon works of Act3D [gervet2023act3d] from 3D policies and upon the works of [urain2023se, chi2023diffusion] from diffusion policies. It uses a tokenized 3D scene representation, similar to [gervet2023act3d], but it is a probabilistic model instead of a deterministic one. It does not sample 3D points and does not infer 3D action maps. It uses diffusion objectives instead of classification or regression objectives used in [gervet2023act3d]. In contrast to [simeonov2023shelving, urain2023se], it uses 3D scene representations instead of 2D images or low-dimensional states. We compare against both 2D diffusion policies and 3D policies in our experiments and show 3D Diffuser Actor greatly outperforms them. We highlight the differences between our model and related models in Figure 3 in the Appendix, and we refer to Figures 4 and 5 for more architectural details of 3D Diffuser Actor and Act3D.
Figure 1: Overview of 3D Diffuser Actor. (a) 3D Diffuser Actor is a conditional diffusion probabilistic model of robot 3D trajectories. At diffusion step $i$, it converts the current noised estimate of the robot’s future action trajectory $\bm{\tau}^{i}$, the posed RGB-D views $\mathbf{o}$, and proprioceptive information $c$ to a set of 3D tokens. It fuses information among these 3D tokens and language instruction tokens $l$ using 3D relative denoising transformers to predict the noise of 3D robot locations $\epsilon^{loc}_{\theta}(\mathbf{o},l,c,\bm{\tau}^{i},i)$ and the noise of 3D robot rotations $\epsilon^{rot}_{\theta}(\mathbf{o},l,c,\bm{\tau}^{i},i)$. (b) During inference, 3D Diffuser Actor iteratively denoises the noised estimate of the robot’s future trajectory. (c) 3D Diffuser Actor works in the real world and captures the multiple behavioural modes in the training demonstrations.
3 Method